### v2

# Learning How to Use Tools, Not Just When: Pattern-Aware Tool-Integrated Reasoning

Abstract
Tool-integrated reasoning (TIR) has become a key approach for improving large reasoning models (LRMs) on complex problems. Prior work has mainly studied when to invoke tools, while overlooking how tools are applied. We identify two common patterns: a calculator-pattern that uses code for direct computation, and an algorithmic-pattern that encodes problems as programs. Misaligned choices often cause failures even when reasoning is sound. We propose a two-stage framework that first builds code competence from both patterns and then aligns pattern selection with teacher preferences. Across challenging math datasets, our pattern-aware method substantially improves both code usage and accuracy—for instance, raising Code@1 on MATH500 from 64.0% to 70.5% and on AIME24 from 26.7% to 50.0%. These gains highlight the effectiveness of a pattern-aware approach for tool-integrated reasoning.
1 Introduction
Tool-integrated reasoning (TIR) has become a powerful paradigm for enhancing large reasoning models (LRMs) wei2025autotir; chen2025r1; yu2025chain. By interacting with external verifiers such as code interpreters, TIR enables models to produce executable reasoning steps, making answers more accurate, faithful, and verifiable zheng2025deepresearcher. Recent progress has shown gains from better timing of tool invocation, multi-round tool calls, and tighter integration of external computation dong2025tool; wei2025autotir.
While effective, a critical question remains: can models reliably decide how to use tools once invoked? We provide an initial exploration of this problem. Models often misalign their tool-use strategies with problem demands, applying code in a mechanical and context-insensitive way that prevents the tool from working effectively. As Figure 1 illustrates, on the factorial problem $1000!\div(800!\times 2!)$ the model used the calculator-pattern and failed due to overflow, while on the “first 10-digit prime in $\pi$” problem it used the algorithmic-pattern and stalled without output. In both cases, switching to the complementary pattern succeeds, indicating that the failures stem from pattern selection rather than a lack of reasoning ability.
To address this challenge, we propose a two-stage learning framework. In the first stage, code competence acquisition, the model learns to generate reliable code from both calculator- and algorithmic-pattern data, without teacher preferences, ensuring stable tool use. In the second stage, pattern preference alignment, teacher signals guide the model to choose the more suitable pattern for each problem. Together, these stages equip LRMs to not only invoke tools effectively but also adapt their strategy of use.
Figure 1: Pattern mismatch in tool-integrated reasoning. Left: in $1000!\div(800!\times 2!)$, an algebraic approach succeeds while direct computation overflows. Right: in finding the first 10-digit prime in $\pi$, a symbolic approach fails from context limits while a scanning approach succeeds. These cases show that success depends on the chosen tool-use pattern.
Our experiments on challenging math benchmarks show substantial and consistent gains. The model’s ability to produce executable code (@1) increases from 64.0% to 70.5% on MATH500, indicating that code competence acquisition markedly stabilizes tool use. At the same time, overall problem-solving accuracy rises from 26.7% to 50.0% on AIME24, demonstrating that pattern preference alignment further improves strategy selection and reduces failures due to mismatched tool use. Together, these results validate that modeling how tools are used—beyond merely invoking them—yields large improvements in both code reliability and end-task performance.
This work makes three contributions. (1) We identify and formalize pattern mismatch in TIR, showing that misaligned tool-use strategies can derail otherwise sound reasoning. (2) We propose a two-stage framework that first builds code competence from both calculator- and algorithmic-pattern data (without teacher preferences) and then aligns pattern choice with teacher signals. (3) We provide empirical evidence on MATH500 and AIME24 that this approach substantially improves code@1 (64.0% $\rightarrow$ 70.5%) and accuracy (26.7% $\rightarrow$ 50.0%), and we illustrate how switching patterns flips failure cases into successes (Fig. 1).
2 Related Works
2.1 Reasoning with LRMs
LRMs have demonstrated better performance when explicitly generate intermediate reasoning stepsxu2025towards. Subsequent work has focused on improving the quality of these reasoning paths, either by introducing diversity through self-consistency wang2022self, refining intermediate steps via process supervision lightman2023let; jiang2025drp, or encouraging verification of intermediate results zelikman2022star; uesato2022solving.
2.2 Tool-Integrated Reasoning
Tool-integrated reasoning has emerged as a powerful paradigm for enhancing large reasoning models (LRMs) wei2025autotir; chen2025r1, enabling executable reasoning steps through external tools such as Python interpreters zheng2025deepresearcher. Early work like PoT chen2022program, PAL gao2023pal, and MathPrompter imani2023mathprompter showed that converting reasoning into code execution or lightweight snippets can substantially improve math performance, while implementations such as ToRA gou2023tora and Qwen2.5-Math-Instruct-TIR shao2024deepseekmath further demonstrated gains from deeper integration. However, prior studies primarily focus on when to call tools, overlooking how tools are applied once invoked. We address this gap by modeling pattern selection in TIR, introducing a pattern-aware framework that aligns tool-use strategies and yields significant improvements in reasoning accuracy.
Figure 2: Two-stage training framework. In the 1st phase, each problem is expanded into both calculator- and algorithmic-style solutions ($2N$ data) to build code competence, and in the 2nd phase, teacher-provided preferences on NN problems guide the student to become pattern-aware.
3 Method
We train a pattern-aware reasoning model in two stages. For stage 1, we builds code competence by supervising the model on both calculator-style and algorithmic-style solutions. Stage 2 performs pattern preference alignment, using Direct preference optimization (DPO) to learn the pattern preference.
3.1 Training set construction
We distinguish two usage patterns in tool-integrated reasoning: the calculator-pattern, where code is used only for direct computation or verification, and the algorithmic-pattern, where problems are expressed as full programs. For each problem, we generate a Chosen Solution with the more suitable pattern and a Counter Solution with the alternative one. This dual construction not only equips the model with competence in both styles of code usage, but also provides the supervision necessary to align its pattern selection with teacher preferences, thereby improving robustness and accuracy in downstream reasoning tasks. And We put the complete version is provided in Section A.
3.2 Code Competence Acquisition
For each problem $x$, we construct two solutions:
  • an algorithmic-style solution $y^{alg}$ that converts the problem to an execuable program;
  • a calculator-style solution $y_{calc}$ that keeps the reasoing path while using python for arithmetic or verification.
We write both solutions with the same ouput shcema (reasoning $\rightarrow$ python block $\rightarrow$ ouputs $\rightarrow$ final answer), which can simplifies parsing and execution. Let $D_{SFT}=\{(x_{i},y_{i}^{alg}),(x_{i},y_{i}^{calc})\}$ be the set of both patterns, we minimize the negative log-likelihood over response tokens:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{SFT}}}\sum_{t=1}^{|y|}\log\pi_{\theta}(y_{t}\mid x,y_{<t}),$
3.3 Pattern Preference Alignment
We use DPO rafailov2023direct to empower our model with the ability to pick the correct pattern to solve math problems. For each problem $x$, we collect two candidate completions $y^{alg},y^{calc}$ and label a winner $y^{+}$ and loser $y^{-}$ using the teacher model. For a pair $(x,y^{+},y^{-})$, we then apply direction preference optimization with a frozen reference policy $\pi_{ref}$ (our SFT model):
$\mathcal{L}_{\text{DPO}}=-\log\sigma\Big(\beta\big[\log\pi_{\theta}(y^{+}\!\mid x)-\log\pi_{\theta}(y^{-}\!\mid x)-\big(\log\pi_{\text{ref}}(y^{+}\!\mid x)-\log\pi_{\text{ref}}(y^{-}\!\mid x)\big)\big]\Big),$
where $\beta$ scales preference strength and $\sigma$ is the logistic function. This increases the likelihood of preferred (pattern-appropriate) solutions relative to dispreferred ones, without training a separate reward model.
4 Experiments
4.1 Dataset
We use google Gemini-2.5-flash-lite as our teacher model, using our designed two pattern prompt to solve the problems in OpenR1-Math-220k openr1math220k dataset. From this corpus, we select the first 10,000 problems and divide them into training and validation sets using a 9:1 ratio. To assess complex mathematical reasoning and generalization, we evaluate on a broad set of out-of-domain benchmarks, including MATH500 math, AIME24 aime2024, and AMC23 amc2024.
4.2 Evaluation Metrics
We report Pass@1 as the primary accuracy metric. For additional insight into reasoning behavior, we also track (i) the proportion of problems where the model’s output contains executable code (Code@1), and (ii) the proportion of problems where such code is both present and leads to the correct final answer (Code+Pass@1).
5 Main Results
Method R1Math GSM8K (OOD) MATH500 (OOD) AIME24 (OOD)
Code@1 Code+Pass@1 Code@1 Code+Pass@1 Code@1 Code+Pass@1 Code@1 Code+Pass@1
R1-Distill-Qwen-1.5B
Base 3.0% 2.0% 4.2% 4.2% 8.0% 6.0% 0.0% 0.0%
+SFT 66.0% 55.5% 72.0% 70.4% 64.0% 58.2% 26.7% 6.6%
+SFT+DPO 72.2% 70.3% 80.1% 77.4% 70.5% 63.2% 50.0% 26.7%
Table 1: Pass@1 and Code+Pass@1 accuracy on R1-Distill-Qwen models across R1Math, GSM8K, MATH500, and AIME24.
According to Table 1, the base model achieves near-zero code usage and accuracy. With SFT, the model begins to leverage Python extensively (e.g., R1Math Code@1 rises to 66%), and correctness follows accordingly. Adding DPO yields further substantial improvements, notably pushing GSM8K Code@1 from 72.0% to 80.1% and AIME24 from 26.7% to 50.0%, while also boosting correctness (e.g., AIME24 Code+Pass@1 from 6.6% to 26.7%). These results highlight that SFT is crucial for enabling tool use, whereas DPO markedly enhances both adoption and accuracy.
6 Conclusion
In this work, we studied two complementary tool-use patterns—algorithmic code generation and calculator-style usage—and showed that many failures arise from mismatched pattern choices rather than reasoning ability. Experiments on GSM8K, MATH500, and AIME24 demonstrate that our pattern-aware approach improves strategy selection and accuracy, offering an effective and lightweight way to enhance tool-integrated reasoning.
Appendix A Complete Prompt
Double Pattern Prompt Use a python interpreter as a tool to solve the following math problem. There are two possible usage patterns: Pattern A: Treat the problem as a coding problem, write a complete solution using Python code. Pattern B: Treat the Python interpreter as a simple calculator, only using it for arithmetic or verification when needed. Your task: 1. Decide which pattern (A or B) is more appropriate for this problem. Call this the Chosen Solution. 2. Provide the Chosen Solution with EXACTLY ONE continuous reasoning paragraph (no lists, no numbering, no bullet points). Then give ONE Python code block, then the code outputs, then the final answer. 3. Provide the Counterfactual Solution using the other pattern with the SAME constraints (one continuous reasoning paragraph, one code block, outputs, final answer). 4. Strict constraints: - Do NOT restate the problem text. - Do NOT include introductions, apologies, or meta comments. - Do NOT duplicate any content across fields. Output format (pure JSON, no backticks, no extra keys):
Appendix B Evaluate Pattern Chosen
To further examine the reliability of Gemini’s pattern selection, we asked GPT-5 to independently select patterns without prior exposure to Gemini’s choices. Across 100 test cases, GPT-5’s selections were consistent with Gemini’s in 98 instances, corresponding to a 98% agreement rate. This high level of concordance provides strong empirical support for the robustness of our approach. This follows LLM as a Judge li2025generation; gao2023human; jiang2024memorization method.